OpenStudy (aryana_maria2323):

Use your graphing calculator to evaluate limit as x goes to infinity of the quantity 1 plus x all raised to the power of 4 divided by x

OpenStudy (aryana_maria2323):

Can you please help? @Abhisar

OpenStudy (astrophysics):

\[\lim_{x \rightarrow \infty} \frac{ (1+x)^4 }{ x }\] this?

OpenStudy (aryana_maria2323):

sorry this is it \[\lim_{x \rightarrow \infty}\left[ 1+x \right]^{\frac{ 4 }{ x }}\]

OpenStudy (astrophysics):

Try using exponentials and logarithms

OpenStudy (aryana_maria2323):

I tried but I still don't understand.

OpenStudy (astrophysics):

Can you show me what you did

OpenStudy (aryana_maria2323):

\[\lim_{x \rightarrow \infty} e ^{\ln \left( 1+x \right)}\]

OpenStudy (aryana_maria2323):

but that is not even close to any of the answers @Astrophysics

OpenStudy (aryana_maria2323):

I am sorry I am having computer problems so I did not see what you guys said. @Loser66 @Astrophysics

OpenStudy (astrophysics):

 continue